# Multi-critic Learning for Whole-body End-effector Twist Tracking

### v1

@@ Abstract
Learning whole-body control for locomotion and arm motions in a single policy has challenges, as the two tasks have conflicting goals. For instance, efficient locomotion typically favors a horizontal base orientation, while end-effector tracking may require the base to tilt to extend reachability. Additionally, current Reinforcement Learning (RL) approaches using a pose-based task specification lack the ability to directly control the end-effector velocity, making smoothly executing trajectories very challenging. To address these limitations, we propose an RL-based framework that allows for dynamic, velocity-aware whole-body end-effector control. Our method introduces a multi-critic actor architecture that decouples the reward signals for locomotion and manipulation, simplifying reward tuning and allowing the policy to resolve task conflicts more effectively. Furthermore, we design a twist-based end-effector task formulation that can track both discrete poses and motion trajectories. We validate our approach through a set of simulation and hardware experiments using a quadruped robot equipped with a robotic arm. The resulting controller can simultaneously walk and move its end-effector and shows emergent whole-body behaviors, where the base assists the arm in extending the workspace, despite a lack of explicit formulations.
@@ 1 Introduction
However, the field of whole-body control using RL has many challenges that are still being addressed. Firstly, locomotion and end-effector control have some conflicting goals. For example, locomotion tends towards a horizontal body orientation for an energy-efficient gait [15, 16, 17], but at the same time, tilting the base can increase the reachability of the arm. Similarly, velocity tracking of the base is far coarser than is needed to achieve precise end-effector control. Whole-body policies that can both locomote and move the arm tend to have low accuracy [7, 1]. Some solutions have attempted to either have separate base and arm controllers [12] or different policies for whole-body arm motion and locomotion [8]. However, these solutions all have limitations, either in requiring switching between controllers, not utilizing the entire potential workspace, or not being able to perform arm movements while in motion. A second challenge in whole-body control with RL is formulating the end-effector tracking task. Existing solutions use various ways to define goal pose tracking [13, 1, 10, 19, 8], or to perform end-effector force tracking [11]. However, trajectory following using a goal pose tracking controller may not always lead to smooth end-effector control since the controller attempts to rigidly track each control point. This is a fundamental issue with such task formulation due to its lack of velocity specification.
To address the first challenge of diverging goals, we propose using a multi-critic actor architecture [20, 21] that decouples the reward signals for locomotion, manipulation, and contact scheduling into individual reward groups. This simplifies the reward tuning, as separate critics allow the policy to resolve task conflicts more effectively. Furthermore, having separate critics seamlessly combines dense tracking rewards with sparse contact schedule rewards. We show that this approach is easier to use and also has desirable emergent behaviors.
As a second contribution, we design a 6D twist-based end-effector task formulation that enables direct control over the end-effector velocity while precisely tracking both discrete poses and continuous motion trajectories. Through comprehensive simulation and hardware experiments, we validate that the twist-based formulation leads to smooth end-effector tracking during stationary operation and during dynamic locomotion of the quadrupedal robot. In summary, our contributions are:
@@ 2 Related Work
Conflicting goals also arise when training whole-body control policies. For example, the reward terms for locomotion tend to favor an overall horizontal base orientation while not enforcing a strict pitch or roll target, allowing for energy-efficient gaits to emerge [15, 16, 17, 18]. These controllers often do not even track small velocity changes. However, when training with end-effector accuracy in mind, a very stiff base is desired, which can also tilt to extend the reachability of the arm. Methods that attempt to balance this tuning problem, such as the ones presented by Fu et al. [7] or Ha et al. [1], usually suffer from accuracy or stability issues, as both approaches have end-effector tracking errors in the tens of centimeters. On the other hand, Portela et al. [8] split the problem into two separate whole-body controllers, one for walking and another for end-effector motions. Their approach can effectively tackle both the accuracy and robustness issues, but requires constant switching between controllers, and the arm can not be moved while the base is walking. In legged-wheeled systems, arm motions are more easily executed, as the presence of wheels reduces the necessity for base movements associated with walking gaits [9, 10]. Our proposed approach tackles this problem using a multi-critic approach [20, 21], enabling the locomotion and end-effector tracking problems to have separate sets of rewards, but one single policy. This emerging approach has been explored for style mixing [20] or locomotion [27, 28]. We showcase further uses and how multi-critic RL training simplifies the tuning problem and emergently leads to desired behaviors.
Another critical aspect of using RL for end-effector tracking is the target and reward formulation for arm movement. Early works focused solely on 3D position tracking [7, 6], which later evolved into full 6D pose tracking, using a variety of representations, from Euler angles [13], to rotation matrices [1, 10, 19], and keypoints [8]. However, for pose-based approaches, smoothly tracking a trajectory can be difficult and requires a high-level policy to monitor the execution [1, 9]. An alternative to pose tracking is force tracking, as demonstrated by Portela et al. [11], which allows the possibility of building an impedance controller on top. In contrast, our approach circumvents these issues by directly performing twist tracking, resulting in simpler reward tuning for the entire problem. Furthermore, the twist tracking formulation provides explicit control over end-effector velocity, enabling the robot to follow complex trajectories.
@@ 3.1 Task definition
Given the observations, the policy $\pi_{\theta}^{T}$ generates an action $\bm{a}_{t}=[\bm{a}_{t}^{arm},~{}\bm{a}_{t}^{robot}]$ where $\bm{a}_{t}^{arm}\in\mathbb{R}^{6}$ and $\bm{a}_{t}^{robot}\in\mathbb{R}^{12}$ are the relative offsets to be applied. Unlike the standard approach[8] of applying actions to predefined default joint positions, we apply these offsets directly to the current joint positions. Assuming a Gaussian action distribution, the actions are normalized with a fixed standard deviation $\sigma_{a}$ and clamped to be within $[-1,1]$. The joint position targets for the next time step $\bm{q}_{t+1}$ are computed as,
@@ 3.2 Command Formulation
End-effector trajectory sampling: During training, we generate manipulation trajectories by initializing the start pose $\bm{T}_{s}$ at the current end-effector location, while the goal pose $\bm{T}_{g}$ is randomly sampled using a two-stage approach. First, we sample positions within a unit sphere centered at the robot’s shoulder. Any samples falling within a predefined cuboid representing the robot’s torso and hip region are rejected to ensure physically feasible goals, which enables smoother training [8]. For orientation determination, we construct a reference orientation where the x-axis aligns with gravity and the z-axis points along the vector from the torso center to the sampled position. This reference orientation is then perturbed with random rotations about each axis, bounded within ±$\textpm\frac{\pi}{4}$ radians. This approach ensures that a wide range of the sampled poses are within the robot’s feasible workspace. Given the start and goal pose, we implement a temporal trajectory formulation that prioritizes continuous tracking rather than terminal goal achievement. We assign a random duration to each trajectory and generate intermediate waypoints to be tracked using linear interpolation in position and spherical linear interpolation in orientation. The intermediate goals are defined as
@@
Foot trajectory: Foot trajectory generation computes the desired foot heights above the ground during locomotion. To achieve this, a gait phase $\phi$ and foot phase offsets $\bm{\theta}=[\theta_{LF},~{}\theta_{RF},~{}\theta_{LH},~{}\theta_{RH}]$ are used to generate a sinusoidal swing foot trajectory pattern [14], where the subscripts denote the four legs: Left Front (LF), Right Front (RF), Left Hind (LH), and Right Hind (RH) . The height of the feet is computed as:
@@
Base command: The base command consists of the linear velocity of the torso ${}_{B}\bm{v}_{b}=[_{B}\bm{v}_{b_{x}},_{B}\bm{v}_{b_{y}}]$ and the angular velocity ${}_{B}\bm{\omega}_{b}$ along the Z-axis wrt. the base frame. During training, this velocity command is periodically resampled during the episode from a uniform distribution within the range $[-0.25,0.25]$.
@@ 3.3 Multi-critic Actor Learning
$L(\theta)=\hat{\mathbb{E}}[min(r_{t}(\theta)\hat{A},~{}clip(r_{t}(\theta),~{}1-\epsilon,~{}1+\epsilon)\hat{A})]$
@@
  • The locomotion group encompasses base velocity tracking rewards and regularization rewards for locomotion actions.
  • The manipulation group includes end-effector position and orientation tracking rewards, along with associated regularization terms.
  • The contact schedule group comprises rewards for tracking foot contact states and corresponding regularization terms.
@@
3.4 Teacher-Student Distillation
Our teacher policy is a 3-layer MLP with hidden layers of size $[512,256,128]$ with ELU activations. The teacher has access to privileged information inaccessible in the real world. To deploy the policy on hardware, a student policy with the same network structure as the teacher is distilled without access to privileged information using supervised learning [16].
@@ 4.1 Hardware Experiments
Tracking performance of different trajectory types: To assess the tracking performance of the policy, we evaluate its ability to track the following trajectory types:
@@
During these experiments, the end effector orientation is actively controlled to point outwards along the vector from the torso center to the end effector while aligning the X-axis of the end-effector frame with gravity vector. In Fig. 3, we show the tracking performance of the policy for all three trajectory types.
Tracking trajectories with varying velocities:
@@
Table 1: Tracking performance in end-effector position error (L2 norm) $\delta\bm{r}$ (m) and velocity error $\delta\dot{\bm{r}}$ (m/s) for robot linear motions XYZ executed at specified distances and speeds $\hat{\bm{v}}_{EE}$ (m/s).
To demonstrate the capability of the controller to execute trajectories at varying velocities, we execute the three trajectory types described above at different commanded velocities. The RMSE end-effector tracking error in position and velocity while tracking trajectories at different speeds is listed in Table 1. As also shown in Fig. 3, the controller achieves good velocity tracking across different velocity ranges. Although velocity tracking accuracy is somewhat lower at slower speeds, particularly for linear trajectories, the policy successfully maintains low position errors throughout the execution. It must be noted that circular and semicircular trajectories being evaluated are not part of the training distribution, since the trajectories in training are always computed between start and goal pose through linear interpolation.
Chicken head control performance while walking:
Figure 4: End-Effector Tracking performance in chicken-head mode: Tracking error in end effector position (red, left axis) and measured base velocity (blue, right axis) of the robot plotted over time. The position error has a mean of 0.0156 m and maximum of 0.0287 m.
To demonstrate the whole-body task coordination capabilities of the controller, we perform chicken head control with the end-effector while walking. The goal of this experiment is to demonstrate the whole-body loco-manipulation capability of the policy by holding a static end-effector pose in the world frame while moving the torso. During this experiment, the robot is commanded a base velocity within the range of [-0.2, 0.2] m/s while the end-effector is commanded twist corrections to hold the end-effector position in the robot’s odometry frame. As shown in Fig. 4, the policy achieves very low end-effector position tracking errors while walking, demonstrating effective task space control during base motion.
4.2 Simulation Experiments
Effect of multi-critic architecture on gait tracking: Multi-critic architectures have previously demonstrated effectiveness in multi-task RL problems. While the agent is trained only with a static walking gait pattern, we observe that the policy generalizes to perform a trotting gait at runtime by simply modifying the commanded foot swing trajectories. The measured foot trajectories for both static walk and trotting gaits are shown in Fig. 5.
@@ 4.2 Simulation Experiments
Comparison with existing methods: To evaluate our proposed approach against existing methods, we conduct comparative experiments using a semicircular end-effector trajectory, both while the robot is stationary and during locomotion. We integrate our arm control with three different locomotion policies: Ma [12], Taka 3-DoF [17], and Taka 6-DoF [30]. To achieve this, we apply the leg actions generated by the respective locomotion policy while executing the arm actions produced by our proposed approach. We compute the tracking end effector tracking errors using the measurements obtained from the state estimator. As shown in Fig. 6, while all combinations achieve satisfactory end-effector tracking, our proposed whole-body locomotion policy demonstrates superior performance with the lowest tracking errors.
@@
Figure 6: Average end-effector tracking errors: $\delta\bm{r}$ (mean Euclidian norm) for position and $\delta\bm{\theta}$ (mean angular error) for orientation, across controllers while standing and walking.

### v2

@@ Abstract
Learning whole-body control for locomotion and arm motions in a single policy has challenges, as the two tasks have conflicting goals. For instance, efficient locomotion typically favors a horizontal base orientation, while end-effector tracking may require the base to tilt to extend reachability. Additionally, current Reinforcement Learning (RL) approaches using a pose-based task specification lack the ability to directly control the end-effector velocity, making smoothly executing trajectories very challenging. To address these limitations, we propose an RL-based framework that allows for dynamic, velocity-aware whole-body end-effector control. Our method introduces a multi-critic actor architecture that decouples the reward signals for locomotion and manipulation, simplifying reward tuning and allowing the policy to resolve task conflicts more effectively. Furthermore, we design a twist-based end-effector task formulation that can track both discrete poses and motion trajectories. We validate our approach through a set of simulation and hardware experiments using a quadruped robot equipped with a robotic arm. The resulting controller can simultaneously walk and move its end-effector and shows emergent whole-body behaviors, where the base assists the arm in extending the workspace, despite a lack of explicit formulations. Videos and supplementary material can be found at multi-critic-locomanipulation.github.io.
@@ 1 Introduction
Whole-body control using RL faces several challenges. Firstly, locomotion and end-effector control have conflicting goals. Locomotion tends towards a horizontal body orientation for energy-efficient gaits [15, 16, 17], but tilting the base can increase the reachability of the arm. Similarly, velocity tracking of the base is far coarser than needed to achieve precise end-effector control. Similarly, the coarse velocity tracking of the base is insufficient for precise end-effector control, leading to low accuracy in unified policies [7, 1]. Solutions using separate controllers [12] or different policies [8] have limitations, such as requiring controller switching, not using the full workspace, or being unable to perform arm movements while in motion. A second challenge is formulating the end-effector tracking task. Current methods use goal pose tracking [13, 1, 10, 19, 8] or force tracking [11]. However, trajectory following using a goal pose tracking controller may not always lead to smooth end-effector control since the controller attempts to rigidly track each control point. This is a fundamental issue with such task formulation due to its lack of velocity specification.
To address the first challenge of diverging goals, we propose using a multi-critic actor architecture [20, 21] that decouples the reward signals for locomotion, manipulation, and contact scheduling into individual reward groups. This simplifies the reward tuning, as separate critics allow the policy to resolve task conflicts more effectively. Furthermore, having separate critics seamlessly combines dense tracking rewards with sparse contact schedule rewards. We show that this approach is easier to use and also has desirable emergent behaviors. As a second contribution, we design a 6D twist-based end-effector task formulation that enables direct control over the end-effector velocity while precisely tracking both discrete poses and continuous motion trajectories. Through comprehensive simulation and hardware experiments, we validate that the twist-based formulation leads to smooth end-effector tracking during stationary operation and during dynamic locomotion of the quadrupedal robot. In summary, our contributions are:
@@ 2 Related Work
Conflicting goals also arise when training whole-body control policies. Locomotion rewards favor horizontal base orientation for energy-efficient gaits [15, 16, 17, 18], while end-effector accuracy requires a stiff, potentially tilted base to extend arm reach. Methods balancing this trade-off [7, 1] suffer from accuracy issues with tracking errors in tens of centimeters. Portela et al. [8] splits walking from end-effector motions but requires controller switching. In legged-wheeled systems, arm motions are more easily executed, as the presence of wheels reduces the necessity for base movements associated with walking gaits [9, 10]. Our proposed approach tackles this problem using a multi-critic approach [20, 21], enabling the locomotion and end-effector tracking problems to have separate sets of rewards, but one single policy. This emerging approach has been explored for style mixing [20] or locomotion [27, 28]. We showcase further uses and how multi-critic RL training simplifies the tuning problem and emergently leads to desired behaviors, benefiting from accuracy and synergy between differing tasks on hardware.
Another critical aspect is target formulation for arm movement. Early works used 3D position tracking [7, 6], evolving to 6D pose tracking with various representations [13, 1, 10, 19, 8]. However, pose-based approaches struggle with smooth trajectory tracking and require high-level monitoring [1, 9]. An alternative to pose tracking is force tracking, as demonstrated by Portela et al. [11], which allows the possibility of building an impedance controller on top. In contrast, our approach circumvents these issues by directly performing twist tracking, resulting in simpler reward tuning for the entire problem. Furthermore, the twist tracking formulation provides explicit control over end-effector velocity, enabling the robot to follow complex trajectories.
@@ 3.1 Task definition
Given the observations, the policy $\pi_{\theta}^{T}$ generates an action $\bm{a}_{t}=[\bm{a}_{t}^{arm},~\bm{a}_{t}^{robot}]$ where $\bm{a}_{t}^{arm}\in\mathbb{R}^{6}$ and $\bm{a}_{t}^{robot}\in\mathbb{R}^{12}$ are the relative offsets to be applied. Unlike the standard approach[8] of applying actions to predefined default joint positions, we apply these offsets directly to the current joint positions. Assuming a Gaussian action distribution, the actions are normalized with a fixed standard deviation $\sigma_{a}$ and clamped to be within [1,1][-1,1]. The joint position targets for the next time step $\bm{q}_{t+1}$ are computed as,
@@ 3.2 Command Formulation
End-effector trajectory sampling: During training, we generate manipulation trajectories by initializing the start pose $\bm{T}_{s}$ at the current end-effector location, while the goal pose $\bm{T}_{g}$ is randomly sampled using a two-stage approach. First, we sample positions within a unit sphere centered at the robot’s shoulder. Any samples falling within a predefined cuboid representing the robot’s torso and hip region are rejected to ensure physically feasible goals, which enables smoother training [8]. For orientation determination, we construct a reference orientation where the x-axis aligns with gravity and the z-axis points along the vector from the torso center to the sampled position. This reference orientation is then perturbed with random rotations about each axis, bounded within $\textpm\frac{\pi}{4}$ radians. This approach ensures that a wide range of the sampled poses are within the robot’s feasible workspace. Given the start and goal pose, we implement a temporal trajectory formulation that prioritizes continuous tracking rather than terminal goal achievement. We assign a random duration to each trajectory and generate intermediate waypoints to be tracked using linear interpolation in position and spherical linear interpolation in orientation. The intermediate goals are defined as
@@
Foot trajectory generation: computes the desired foot heights above the ground during locomotion. A gait phase $\phi$ and foot phase offsets $\bm{\theta}=[\theta_{LF},~\theta_{RF},~\theta_{LH},~\theta_{RH}]$ are used to generate a sinusoidal swing foot trajectory pattern [14], where the subscripts denote the four legs: Left Front (LF), Right Front (RF), Left Hind (LH), and Right Hind (RH) . The height of the feet is computed as:
@@
The base command: consists of the linear velocity of the torso $_{B}\bm{v}_{b}=[_{B}\bm{v}_{b_{x}},_{B}\bm{v}_{b_{y}}]$ and the angular velocity $_{B}\bm{\omega}_{b}$ along the Z-axis wrt. the base frame. During training, this velocity command is periodically resampled during the episode from a uniform distribution within the range [0.25,0.25][-0.25,0.25].
@@ 3.3 Multi-critic Actor Learning
[…]
@@
  • Locomotion: base velocity tracking rewards and regularization terms.
  • Manipulation: end-effector pose tracking rewards and regularization terms.
  • Contact schedule: rewards and regularization terms for tracking foot contact states.
@@ 4 Results
Tracking performance of different trajectory types on hardware: To assess the tracking performance of the policy, we evaluate its ability to track the following trajectory types:
@@
During these experiments, the end effector orientation is actively controlled to point outwards along the vector from the torso to the end effector while aligning the X-axis of the end-effector frame with gravity vector. In Fig. 3, we show the tracking performance for all three trajectory types.
Tracking trajectories with varying velocities on hardware:
@@
Table 1: Tracking performance in end-effector position error (L2 norm) $\delta\bm{r}$ (m) and velocity error $\delta\dot{\bm{r}}$ (m/s) for robot linear motions XYZ executed on hardware at specified distances and speeds $\hat{\bm{v}}_{EE}$ (m/s).
To demonstrate the capability of the controller to execute trajectories at varying velocities, we execute the three trajectory types described above at different commanded velocities. The RMSE end-effector tracking error in position and velocity while tracking trajectories at different speeds is listed in Table 1. As also shown in Fig. 3, the controller achieves good velocity tracking across different velocity ranges. Although the accuracy is lower at slower speeds, particularly for linear trajectories, the policy maintains low position errors throughout the execution. It must be noted that circular and semicircular trajectories being evaluated are not part of the training distribution, since the trajectories in training are always computed between start and goal pose through linear interpolation.
Effect of multi-critic architecture on gait tracking in simulation: Multi-critic architectures have previously demonstrated effectiveness in multi-task RL problems. While the agent is trained only with a static walking gait pattern, we observe that the policy generalizes to perform a trotting gait at runtime by simply modifying the commanded foot swing trajectories. The measured foot trajectories for both static walk and trotting gaits are shown in Fig. 4.
@@
Comparison with existing methods in simulation: To evaluate our proposed approach against existing methods, we conduct comparative experiments using a semicircular end-effector trajectory, both while the robot is stationary and during locomotion. We integrate our arm control with three different locomotion policies: Ma [12], Taka 3-DoF [17], and Taka 6-DoF [30].
@@
Table 2: Average end-effector tracking errors: $\delta\bm{r}$ (mean Euclidian norm) for position and $\delta\bm{\theta}$ (mean angular error) for orientation in simulation.
Figure 5: Workspace sweep trajectory wrt. base used for evaluating tracking performance in simulation.
To achieve this, we apply the leg actions generated by the respective locomotion policy while executing the arm actions produced by our proposed approach. We compute the tracking end effector tracking errors using the measurements obtained from the state estimator. As shown in Table 2, while all combinations achieve satisfactory end-effector tracking, our proposed whole-body locomotion policy demonstrates superior performance with the lowest tracking errors.
@@
In addition to the linear and circular trajectories introduced above, we introduce a workspace sweep trajectory that extends the semicircular trajectory by sweeping across various radii and heights. As in Fig. 5, our proposed approach demonstrates improved end-effector tracking performance across the evaluated workspace compared to a pose-based alternative. In Table 3, we present an extended comparison of end-effector tracking performance for linear, circular, and workspace sweep trajectories, and provide results also against a whole-body MPC controller [2]. Our proposed approach consistently outperforms all other methods across all trajectory types and both standing and walking scenarios, achieving the lowest position and orientation errors as well as velocity tracking errors, demonstrating whole-body loco-manipulation capabilities.
Type Controller Standing Walking
$\delta\bm{r}$ [m] $\delta\bm{\theta}$ [] $\delta\bm{\dot{r}}$ [m/s] $\delta\bm{\dot{\theta}}$ [rad/s] $\delta\bm{r}$ [m] $\delta\bm{\theta}$ [] $\delta\bm{\dot{r}}$ [m/s] $\delta\bm{\dot{\theta}}$ [rad/s]
Linear Multi-critic (Ours) 0.0095 1.6084 0.0568 0.1756 0.0093 1.9926 0.0559 0.2176
Taka 3-DoF [17] 0.0474 10.1269 0.2934 1.1056 0.0408 16.5597 0.2514 1.8080
Taka 6-DoF [30] 0.0271 8.2354 0.1648 0.8983 0.0370 9.9586 0.2285 1.0868
Ma [12] 0.0232 8.0863 0.1411 0.8823 0.0573 14.2652 0.3562 1.5563
Portela [8] 0.0240 1.7285 0.1421 0.7562 - - - -
Whole-body MPC 0.0207 3.9522 0.1291 0.4311 0.0195 2.8417 0.1221 0.3100
Circular Multi-critic (Ours) 0.0080 0.9338 0.0506 0.1018 0.0137 1.5921 0.0865 0.1736
Taka 3-DoF [17] 0.0446 12.3341 0.2774 1.3449 0.0259 11.8966 0.1611 1.2986
Taka 6-DoF [30] 0.0234 9.7642 0.1447 1.0658 0.0460 40.1593 0.2865 4.3794
Ma [12] 0.0269 36.6286 0.1663 3.9952 0.0425 7.8882 0.2653 0.8600
Portela [8] 0.0399 2.7328 0.2358 1.1981 - - - -
Whole-body MPC 0.0239 3.2007 0.1497 0.3491 0.0327 4.2238 0.2044 0.4607
Workspace sweep Multi-critic (Ours) 0.0233 3.9257 0.1477 0.4323 0.0408 8.1143 0.2520 0.8876
Taka 3-DoF [17] 0.1415 13.6202 0.8812 1.4878 0.0858 9.2738 0.5329 1.0137
Taka 6-DoF [30] 0.0821 7.6319 0.5094 0.5094 0.0900 9.8273 0.5591 1.0736
Ma [12] 0.0906 8.6375 0.5626 0.9454 0.1514 16.8287 0.9439 1.8366
Portela [8] 0.0486 2.1633 0.2804 0.9737 - - - -
Whole-body MPC 0.3244 38.2293 2.0277 4.1702 0.3346 39.5757 2.0915 4.3170
Table 3: Extended comparison with trajectory and velocity tracking errors in simulation.
Figure 6: Reward sensitivity analysis of single-critic vs multi-critic frameworks: Multi-critic policy is robust against reward scaling across categories whereas the single-critic policy only learns to perform the objective with higher reward scale.
Reward Sensitivity Analysis - Single-Critic vs Multi-Critic Learning: Since our approach splits rewards categorically into locomotion, manipulation, etc., we multiply all rewards within each category by a fixed multiplicative factor and evaluate the resulting policy performance for reward sensitivity analysis. Specifically, we conduct experiments scaling locomotion and manipulation rewards by factors of 5 and 10. As shown in Fig. 6, the multi-critic approach remains insensitive to these weighting factors, while the single-critic approach only learns to perform either locomotion or manipulation and not to perform both simultaneously. When scaling the locomotion reward in the single-critic approach, we observe improved locomotion performance but deteriorated end-effector tracking. In contrast, scaling the manipulation reward improves the end-effector tracking performance but degrades locomotion quality. These experiments demonstrate that the proposed multi-critic approach is robust to variations in relative weightage between reward categories, making it effective for combining multiple learning objectives without careful reward tuning.